Draw the shear and moment diagrams. 2.1 kN/m A C B - 5 m - 2 m
Answer #1

FrWhe! CD 2.1 kN/m A .C sm As 2ml FB RA Efy=0 RA+RB – **7 2.1 RATRB = 7:35. KN 0 ग EMA=0 IX 7 X 2.1 x RB X5 =0 2 [R8 = 6.86 KNSFB SFA = 0.4g. On 15 (0)2 = 0.49 kn 0.49-0.15 (5)2 = -3.26 kN cil portion BCA (5<x< 7m) SFxa - 1 x x x W x + RB s fra 0.49 -BMmap, = 0.49(1.807) 0.49(1.807) -0.05 (1.80733 0:59 KN-m 1 3 + 34.3 (ii) portions BC + (56257 m) Mx = RAx- 1 x x x W x 2 + R